# Multi-LLM Text Summarization

### v1

@@ 1 Introduction
Large language models (LLMs) have been shown to have the potential to produce high-quality summaries Chowdhery et al. (2022); Zhang et al. (2023); Goyal et al. (2023); Pu et al. (2023b). However, despite the remarkable progress in LLM-based summarization, limitations still exist for long documents where useful information is abundant but sparsely distributed throughout the text. Research by Liu et al. (2023) highlights that a naive application of LLMs may overlook critical details or fail to grasp the holistic meaning of a document, indicating the need for more refined methods.
@@
  • We find that multi-LLM text summarization often performs better than using a single LLM for summarization.
@@ 2.1 Summarization
While these studies validate the potential of LLMs in summarizing well-structured and relatively short texts, they fail to address the unique challenges of long-document summarization, where inputs lack clear structural cues and exhibit greater complexity. Research focusing on long-text summarization, such as Keswani et al. (2024), employed semantic clustering and multi-stage summarization with LLaMA2 to manage lengthy inputs. However, such approaches often rely on predefined hierarchical processing strategies that may oversimplify the nuanced relationships within the text. Moreover, as Liu et al. (2023) noted, LLMs tend to neglect content from the middle sections of long documents, resulting in incomplete or unbalanced summaries.
These limitations point to the need for a novel approach capable of addressing the inherent challenges of long-document summarization. Our work builds upon this foundation by proposing a multi-LLM framework designed to overcome these shortcomings through information exchange and collaborative synthesis, which can better capture the diversity and complexity of long texts.
@@ 2.2 Multi-LLM
Although these studies reveal the potential of multi-LLM approaches, their focus remains on structured reasoning tasks, such as question answering and fact-checking. They have not been adequately explored in the context of long-text summarization, where the challenges include synthesizing distributed information, addressing content imbalances, and preserving the coherence of summaries across extended texts.
We hope to bridge this gap by adapting multi-LLM frameworks to the domain of long-document summarization, addressing limitations of both single LLM and traditional hierarchical techniques, and positioning multi-LLM summarization as a promising solution for long-form content.
@@ 3 Multi-LLM Summarization Framework
In this work, we propose a novel multi-LLM summarization framework that leverages multiple large language models to enhance summarization quality of long document input. Through the distribution of generation and evaluation of candidate summaries across multiple models, our framework aims to provide better summaries than single LLM methods, leveraging expertise from different models. We present two interaction topologies, centralized and decentralized, to guide the collaboration, evaluation, and refinement of summaries between LLMs. Visually these two methods can be represented at a high level in Figure 1. Our approach tackles long text document input, which can span to tens of thousands of words and as such usually exceeds the context window of most standard LLMs. To handle this, we establish a two stage process that involves chunking the source document, independently summarizing each chunk of the source document, and then applying a second round of chunking and summarization on the concatenated intermediate results. Throughout both these stages, both frameworks allow multiple LLMs to collaborate and converge on a single final high quality summary of the entire original reference document. Table 1 provides an overview of our framework’s four main variations.
@@ 4.1.1 Generation Phase
This initial generation stage corresponds to lines 57 of Algorithm 1. Conceptually, each model contributes its unique perspective, leading to a diverse pool of candidate summaries, which is important for robust summary selection in the following evaluation phase.
@@ 4.3 Analysis of Complexity
7:         Send $S_{j}^{(i)}$ to agent $C$
8:     Let $\mathcal{S}_{i}=\{S_{1}^{(i)},S_{2}^{(i)},\ldots,S_{k}^{(i)}\}$
9:     $E^{(i)}=C(P_{ec},\mathcal{S}_{i})$
10:     $\bm{\mathrm{r}}=\textsc{AggrResults}(E^{(i)})$
11:     $j\leftarrow\operatorname*{argmax}_{M_{j}\in\mathcal{M}}r_{j}$
12:     Set $S^{*}\leftarrow S_{j}^{(i)}$
13:     if $\textsc{Converged}(\bm{\mathrm{r}})$ then return $S^{*}$
14:     Set $P$ to prompt in Figure 3.
@@ 5.3 Analysis of Complexity
7:         Send $S_{j}^{(i)}$ to $\mathcal{M}$ models
8:     Let $\mathcal{S}_{i}=\{S_{1}^{(i)},S_{2}^{(i)},\ldots,S_{k}^{(i)}\}$
9:     for each model $M_{j}\in\mathcal{M}$  do
10:         $E_{j}^{(i)}=M_{j}(P_{e},S_{1}^{(i)},\ldots,S_{k}^{(i)})$
11:         Send $E_{j}^{(i)}$ to other $\mathcal{M}$ models
12:     Set $\mathcal{E}_{i}=\{E_{1}^{(i)},E_{2}^{(i)},\ldots,E_{k}^{(i)}\}$
13:     $\bm{\mathrm{r}}=\textsc{AggrResults}(E_{1}^{(i)},\ldots,E_{k}^{(i)})$
14:     $j\leftarrow\operatorname*{argmax}_{M_{j}\in\mathcal{M}}r_{j}$
15:     Set $S^{*}\leftarrow S_{j}^{(i)}$
16:     if $\textsc{Converged}(\bm{\mathrm{r}})$ then return $S^{*}$
17:     Set $P$ to prompt in Figure 3.
@@ 6.2 Main Results
Decentralized 3 Rounds GPT-3.5 & GPT-4o mini 0.328 0.167 0.217 0.030
@@
Centralized 3 Rounds GPT-3.5 & GPT-4o mini 0.312 0.163 0.199 0.029
@@ Appendix A Detailed Experimental Setup
Evaluation Metrics: We assess the quality of LLM-generated summaries using ROUGE-1, ROUGE-L, BLEU-1, and BLEU-4 metrics. ROUGE scores emphasizes recall while BLEU scores emphasize precision. Baselines: For comparison with our multi-LLM approach, unless otherwise mentioned, we leverage GPT-3.5, GPT-4o, GPT-4o mini, and LLaMA3-8B as baselines. For these models, we perform the same chunking across all models, and the summarization prompt is identical to that in the first round of the multi-LLM summarization process (Figure 2). Unless otherwise mentioned, all models use 4K-char chunk-size, and the final summary for each document is concatenation of the generated summaries for each chunk in that document. Finally, unless otherwise mentioned, we set $W=160$ for all models.
@@ C.4 Specialized Prompting
ArXiv GovReport
1 round max 0.339 0.180 0.224 0.043 0.468 0.190 0.477 0.112
1 round max 0.333 0.173 0.219 0.036 0.479 0.197 0.485 0.121
1 round max 0.338 0.175 0.236 0.040 0.469 0.181 0.486 0.104
1 round max 0.355 0.181 0.251 0.049 0.482 0.185 0.494 0.115

### v2

@@ 1 Introduction
Large language models (LLMs) have been shown to have the potential to produce high-quality summaries Chowdhery et al. (2022); Zhang et al. (2023); Goyal et al. (2023); Pu et al. (2023b). However, despite the remarkable progress in LLM-based summarization, limitations still exist for documents where useful information may be sparsely distributed throughout the text. Research by Liu et al. (2023) highlights that a naive application of LLMs may overlook critical details or fail to grasp the holistic meaning of a document, indicating the need for more refined methods.
@@
  • We find that multi-LLM text summarization often performs better than using a single LLM for summarization, and we show that the best performing method in the framework aligns with human judgments.
@@ 2.1 Summarization
While these studies validate the potential of LLMs in summarizing well-structured texts, they may falter for inputs lacking clear structural cues and exhibiting greater complexity. Research focusing on longer text summarization, such as Keswani et al. (2024), employed semantic clustering and multi-stage summarization with LLaMA2 to manage lengthy inputs. However, such approaches often rely on predefined hierarchical processing strategies that may oversimplify the nuanced relationships within the text. Moreover, as Liu et al. (2023) noted, LLMs tend to neglect content from the middle sections of longer documents, resulting in incomplete or unbalanced summaries.
Our work aims to improve performance for both long and short text summarization, and it builds upon aforementioned foundations by proposing a multi-LLM framework designed to overcome these shortcomings through information exchange and collaborative synthesis.
@@ 2.2 Multi-LLM
Although these studies reveal the potential of multi-LLM approaches, their focus remains on structured reasoning tasks, such as question answering and fact-checking. They have not been adequately explored in the context of synthesizing distributed information, addressing content imbalances, and preserving the coherence of summaries across extended texts.
We hope to bridge this gap by adapting multi-LLM frameworks to the domain of document summarization, addressing limitations of both single LLM and traditional hierarchical techniques, and positioning multi-LLM summarization as a promising solution.
@@ 3 Multi-LLM Summarization Framework
In this work, we propose a novel multi-LLM summarization framework that leverages multiple large language models to enhance summarization quality of long document input. Through the distribution of generation and evaluation of candidate summaries across multiple models, our framework aims to provide better summaries than single LLM methods, leveraging expertise from different models. We present two interaction topologies, centralized and decentralized, to guide the collaboration, evaluation, and refinement of summaries between LLMs. Visually these two methods can be represented at a high level in Figure 1. In the datasets we test, articles are typically tens of thousands of words long and exceed the context window of most standard LLMs. To handle this, we establish a two stage process that involves chunking the source document, independently summarizing each chunk of the source document, and then applying a second round of chunking and summarization on the concatenated intermediate results. Throughout both these stages, both frameworks allow multiple LLMs to collaborate and converge on a single final high quality summary of the entire original reference document. Table 1 provides an overview of our framework’s four main variations.
@@ 4.1.1 Generation Phase
This initial generation stage corresponds to line 6 of Algorithm 1. Conceptually, each model contributes its unique perspective, leading to a diverse pool of candidate summaries, which is important for robust summary selection in the following evaluation phase.
@@ 4.3 Analysis of Complexity
7:     Let $\mathcal{S}_{i}=\{S_{1}^{(i)},S_{2}^{(i)},\ldots,S_{k}^{(i)}\}$
8:     $E^{(i)}=C(P_{ec},\mathcal{S}_{i})$
9:     $\bm{\mathrm{r}}=\textsc{AggrResults}(E^{(i)})$
10:     $j\leftarrow\operatorname*{argmax}_{M_{j}\in\mathcal{M}}r_{j}$
11:     Set $S^{*}\leftarrow S_{j}^{(i)}$
12:     if $\textsc{Converged}(\bm{\mathrm{r}})$ then return $S^{*}$
13:     Set $P$ to prompt in Figure 3.
@@ 5.3 Analysis of Complexity
7:     Let $\mathcal{S}_{i}=\{S_{1}^{(i)},S_{2}^{(i)},\ldots,S_{k}^{(i)}\}$
8:     for each model $M_{j}\in\mathcal{M}$  do
9:         $E_{j}^{(i)}=M_{j}(P_{e},S_{1}^{(i)},\ldots,S_{k}^{(i)})$
10:     Set $\mathcal{E}_{i}=\{E_{1}^{(i)},E_{2}^{(i)},\ldots,E_{k}^{(i)}\}$
11:     $\bm{\mathrm{r}}=\textsc{AggrResults}(E_{1}^{(i)},\ldots,E_{k}^{(i)})$
12:     $j\leftarrow\operatorname*{argmax}_{M_{j}\in\mathcal{M}}r_{j}$
13:     Set $S^{*}\leftarrow S_{j}^{(i)}$
14:     if $\textsc{Converged}(\bm{\mathrm{r}})$ then return $S^{*}$
15:     Set $P$ to prompt in Figure 3.
@@ 6.2 Main Results
Decentralized 3 Rounds GPT-3.5 & GPT-4o mini 0.313 0.163 0.200 0.029
@@
Centralized 3 Rounds GPT-3.5 & GPT-4o mini 0.329 0.168 0.217 0.031
@@ 6.4 Cost Analysis
6.5 Human evaluation
In addition to the ablation studies, we perform human evaluation of summaries similar to the last step of the multi-LLM framework, i.e. the evaluation phase (Section 4.2.2). The human raters are prompted with two sets of summaries from the generation phase (Section 4.1.1), and are instructed to evaluate these two sets of summaries for Coherence, Conciseness, and Fluency on 5-point Likert scales Conroy and Dang (2008) with rating guidelines for each possible score (Figure 10). The summaries are randomized and anonymized to reduce bias attributable to knowledge of authorship. We drop the Relevance criteria since no original text is provided to the human raters due to length.
We obtain 420 ratings from 7 raters, and find that humans generally prefer summaries produced by GPT-4o mini, which aligns with preferences by our multi-LLM framework. Human preferences also align with machine preferences for all three evaluation criteria to some degree. Conciseness has the highest agreement with multi-LLM evaluations ($\kappa=0.6$). More details for human evaluations can be found in Appendix D.
@@ Appendix A Detailed Experimental Setup
Evaluation Metrics: We assess the quality of LLM-generated summaries using ROUGE-1, ROUGE-L, BLEU-1, and BLEU-4 metrics. ROUGE scores emphasizes recall while BLEU scores emphasize precision.
Baselines: For comparison with our multi-LLM approach, unless otherwise mentioned, we leverage GPT-3.5, GPT-4o, GPT-4o mini, and LLaMA3-8B as baselines. For these models, we perform the same chunking across all models, and the summarization prompt is identical to that in the first round of the multi-LLM summarization process (Figure 2). Unless otherwise mentioned, all models use 4K-char chunk-size, and the final summary for each document is concatenation of the generated summaries for each chunk in that document.
Finally, unless otherwise mentioned, we set $W=160$ for all models.
@@ C.4 Specialized Prompting
ArXiv GovReport
1 round max 0.339 0.180 0.224 0.043 0.468 0.190 0.477 0.112
1 round max 0.333 0.173 0.219 0.036 0.479 0.197 0.485 0.121
1 round max 0.338 0.175 0.236 0.040 0.469 0.181 0.486 0.104
1 round max 0.355 0.181 0.251 0.049 0.482 0.185 0.494 0.115
@@
Appendix D Human Evaluation
In the human evaluation, we select the first 10 pairs of summaries generated before the final evaluation step by the decentralized, one-round maximum framework (the best-performing setup for ArXiv; see Table 2), and prompt human raters to rate each summary according to Coherence, Conciseness, and Fluency metrics, each represented by a 5-point Likert scale. The goal is to compare human rater preferences with preferences of the LLM performing the final evaluation and therefore producing the final result of the multi-LLM pipeline. Each summary pair consists of texts generated by GPT-3.5 and GPT-4o mini randomized in order of presentation and anonymized such that the raters do not know which model produced which summaries. We do not prompt raters with the corresponding original text as in the multi-LLM method due to the original text’s length and technicality, and therefore we remove the Relevance criterion used in Conroy and Dang (2008). For the remaining criteria we provide guidelines for each possible point value to improve reproducibility. Instructions provided to human raters and rating guidelines can be found in Figure 10.
At the end of human evaluation, we collect 420 ratings from 7 raters (6 men, 1 woman; ages 23-31; 4 East Asian, 2 South Asian, 1 White). [^2] Raters are authors of this paper and close associates who have consented to submitting evaluations, and no rater has prior knowledge of authors of the set of summaries being evaluated. To determine preferences for the human raters, we average the rating for each model and each summary in each evaluation criterion (Table 9). Thus, for each summary and for each criterion we obtain two averaged scores, one for GPT-3.5 and GPT-4o mini. We then determine the human preference by choosing the model with the higher score, and if the scores are the same, we fallback on the default choice GPT-3.5, consistent with the fallback default in the evaluation step in our multi-LLM framework (Table 10). We note that in all three criteria our framework show some agreement with the human raters, as measured by Cohen’s kappa. For conciseness, we observe an agreement of $\kappa=0.6$.
Summary Coherence Conciseness Fluency Averaged
Average scores
by raters for: GPT-3.5 GPT-4o mini GPT-3.5 GPT-4o mini GPT-3.5 GPT-4o mini GPT-3.5 GPT-4o mini
1 3.57 3.57 3.85 3.57 4.42 4.28 3.95 3.80
2 4.28 4.00 4.42 3.85 4.85 4.85 4.52 4.23
3 3.42 4.57 3.57 4.42 3.85 4.57 3.61 4.52
4 3.71 4.57 3.42 4.57 4.28 4.85 3.80 4.66
5 3.71 4.14 3.71 3.85 4.00 4.14 3.80 4.04
6 4.00 4.71 3.57 4.14 4.57 4.42 4.04 4.42
7 4.00 4.71 4.00 4.00 4.28 4.71 4.09 4.47
8 4.00 4.57 4.28 4.28 4.57 4.57 4.28 4.47
9 4.00 4.42 3.85 4.14 4.00 4.42 3.95 4.33
10 4.14 3.85 4.42 4.00 4.71 4.57 4.42 4.14
Table 9: Averaged scores (out of 5) given by human raters for each evaluation criterion for the first 10 summaries from the ArXiv dataset. The raters are asked to rate each summary on coherence, conciseness, and fluency on a 5-point Likert scale. We additionally show the score averaged from the scores for the three criteria. We bold the higher average score for each criterion, and underline the choice of the human raters between GPT-3.5 and GPT-4o mini summaries. When two summaries in a particular criterion have the same average score, we fallback on the default choice GPT-3.5, consistent with the evaluation step in our multi-LLM framework.
Human raters Multi-LLM (Ours)
Summary Coherence Conciseness Fluency Averaged
1 GPT-3.5 GPT-3.5 GPT-3.5 GPT-3.5 GPT-3.5
2 GPT-3.5 GPT-3.5 GPT-3.5 GPT-3.5 GPT-4o mini
3 GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini
4 GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini GPT-3.5
5 GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini
6 GPT-4o mini GPT-4o mini GPT-3.5 GPT-4o mini GPT-4o mini
7 GPT-4o mini GPT-3.5 GPT-4o mini GPT-4o mini GPT-3.5
8 GPT-4o mini GPT-3.5 GPT-3.5 GPT-4o mini GPT-3.5
9 GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini GPT-4o mini
10 GPT-3.5 GPT-3.5 GPT-3.5 GPT-3.5 GPT-3.5
Cohen’s $\kappa$ 0.2 0.6 0.1 0.2
Table 10: Human choices obtained from choosing the model with the higher average score, done separately for each evaluating criterion. At the right-most column we show the choices made in the final evaluation step by our multi-LLM framework, specifically the decentralized one-round-max setup with GPT-3.5 as the evaluator. At the bottom row we show the inter-rater agreement (as measured by Cohen’s kappa) between the human choices and the machine choices for each criterion
[…]
4pt7pt
Read the two summaries (column C and H) and grade them on the following aspects Give a grade from 1 - 5 for each category, 1 being the lowest score and 5 being the highest score, as described below. For summary 1, fill in columns E-G; for summary 2, fill in columns J-L.
Coherence
Evaluate the logical flow of sentences and organization of information within the summary.
1: Ideas do not logically follow from one to another, even if they are relevant.
2: There are some phrases to connect one idea to another.
3: Main ideas connect logically with each other. A few arguments/points are out of order.
4: Most ideas follow logically from one to the next. There may be minor points or words/phrases that seem out of place.
5: All ideas, even if irrelevant, are well connected and organized in a clear overall structure.
Conciseness
Penalize summaries that are overly verbose.
1: Ideas are repeated multiple times, or are described in excessive detail or are verbose, even if ideas are logically organized or are relevant.
2: Ideas/arguments/phrases are sometimes repeated. Most ideas are described in too much detail, or, the text is generally verbose with jargon.
3: Ideas are generally described in appropriate detail. Parts of texts may be unnecessarily verbose or use unnecessary jargon.
4: Most ideas are described in appropriate detail. There may be occasional verbosity.
5: All ideas are described with appropriately complex or simple sentences.
Fluency
Evaluate the grammatical correctness of the generated summary.
Check for any awkward phrasing or grammatical errors that might hinder comprehension. 1: Text is grammatically incorrect or very difficult to understand. There are incomplete sentences or incorrectly used punctuations.
2: The text is generally difficult to understand, but some sections convey meaningful ideas.
3: The text is generally grammatically correct. Some sentences may have incorrect grammar.
4: The text is grammatically correct and sentences have understandable structure. There are occasional incorrectly phrased sentences.
[…]
Figure 10: Screenshot of the interface (with scores already filled in) and instructions given to raters for human evaluation. Instructions include the column indices (not shown in the screenshot) for easier reference. Each summary is rated according to the criteria listed in the instructions (i.e. Coherence, Conciseness, and Fluency). We provide guidelines for each criterion and each possible score for that criterion.